A boy traveled 100 km north from his house, traveled 70 km west and finally 60 km north. What is his distance from his house?
Answer by TimothyLamb(4379)

use vector addition:
v(0,100) + v(-70,0) + v(0,60) = v(-70,160)
---
d = vector length = sqrt( xx + yy )
d = sqrt( (-70)*(-70) + 160*160 )
d = sqrt( 4900 + 25600 )
d = sqrt( 30500 )
d = 174.642491966
---
answer:
distance from his house = 174.64 km
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php